OpenStudy (anonymous):

I have to find the exact answer for this. \[\sec \left( \tan^{-1} 5 \right)\]

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of a triangle with an angle for which \(\tan(\theta)=5\) all you are missing is the hypotenuse, which you get via pythagoras

OpenStudy (anonymous):

\[h=\sqrt{1^2+5^2}=\sqrt{26}\] 

OpenStudy (anonymous):

since secant is hypotenuse over adjacent, you get \[\sec(\theta)=\frac{\sqrt{26}}{1}=\sqrt{26}\]

OpenStudy (anonymous):

if you want, redo this with \[\sec(\tan^{-1}(x))\] and get a formula with an \(x\) instead of the number \(5\)

OpenStudy (anonymous):

Wow. So simple I couldn't see it. I was lost because I didn't know where tan 5 was on the unit circle. Geez. Thanks for opening my eyes,

OpenStudy (anonymous):

yw easy when you know what to do right? that is, viewed the right way. like being asked "if the tangent is 5, what is the secant?"